Two small discs of masses m1 and m2 interconnected by a weightless spring rest on a smooth horizontal plane. The discs are set in motion with initial velocities v1 and v2 whose directions are mutually perpendicular and lie in a horizontal plane. Find the total energy E of this system in the frame of the centre of inertia
1
Expert's answer
2020-10-15T02:31:22-0400


Initially "U=\tilde{U}=0"


so "\tilde{E}=\tilde{T}"


"\tilde{T}=\frac{1}{2} \mu v_{rel}^2=\frac{1}{2}\mu|\vec{v_1}-\vec{v_2}|^2"


Therefore from the equation;

"\tilde{T}=\frac{1}{2} \mu |\vec{v_1}-\vec{v_2}|"

"\vec{v_1}-\vec{v_2}"


Hence;

"\tilde{T}=\frac{1}{2}\frac{m_1m_2}{m_1+m_2}(v^2_1+v^2_2)"

where "\mu=\frac{m_1m_2}{m_1+m_2}" refers to reduced mass